An investment of $2000 earns an annual interest rate of 4%.What will be the value of the investment after 5 years?
Answer by stanbon(75887)

An investment of $2000 earns an annual interest rate of 4%.What will be the value of the investment after 5 years?
----------------
If it is simple interest V(5) = (2000 +2000*0.04*5) = (2000)(1+0.04*5)
= #2400
--------------
If it is compound interest V(5)= 2000(1+0.04)^5 = $2433.00
=================
Cheers,
Stan H.